OpenStudy (anonymous):

Simplify this expression (x^-4)^-5(x^-4)^2/(x^3)^7

OpenStudy (anonymous):

well when a exponent with a power is put to another power, the exponents multiply

OpenStudy (anonymous):

so it becomes \[\frac{ x ^{-4\times-5}x ^{-4\times2} }{ x ^{3\times7} }\]

OpenStudy (anonymous):

then that becomes \[\frac{ x ^{20}x ^{-8} }{ x ^{21} }\]

OpenStudy (anonymous):

then now variables that have exponents and are the same variable, the exponents are added so it becomes \[\frac{ x ^{12} }{ x ^{21} }\]

OpenStudy (anonymous):

when divided, the expononets are subtracted \[x^{-9}\]

OpenStudy (anonymous):

x^-9 since that in almost unreadable

OpenStudy (anonymous):

if you want it as a positive exponent the number has to be inverted to it becomes 1/x^9